OpenStudy (anonymous):

Can someone help me with this. A mixture of 30 lbs. of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of each kind were used to make the mixture? ___ lbs. of $1.50 chocolates and ____ lbs. of 90¢ chocolate

OpenStudy (baldymcgee6):

okay, so the bag of 30 lb candy will cost how much total?

OpenStudy (anonymous):

1.10

OpenStudy (baldymcgee6):

the candy is worth 1.10 per lb, we have 30lbs of it, so therefore, 30lb * $1.10 per lb = $33.00

OpenStudy (baldymcgee6):

Does that part make sense?

OpenStudy (anonymous):

15x+33=270-9x

OpenStudy (anonymous):

yes it does

OpenStudy (baldymcgee6):

lets try it this way, x + y = 30 lb $1.5x + $0.90y = $33.00

OpenStudy (baldymcgee6):

Does that part now make sense?

OpenStudy (anonymous):

yes

OpenStudy (baldymcgee6):

Do you know how to solve it?

OpenStudy (anonymous):

are we solving for x

OpenStudy (baldymcgee6):

Well, we have to equations, we will use the method of sub substitution to solve them. Since we have two unknowns (x and y), we need two equations to solve the system. x + y = 30 lb Solve for x x = 30 - y Now we know what x is in terms of y, so lets sub that into our second equation. $1.5y + $0.90x = $33.00 $1.5y + $0.90(30 - y) = $33.00 Now we must solve for y